# Studying the Effects of Collaboration in Interactive Theme Discovery Systems

### v1

@@ 5 Evaluation
In this section, we use both descriptive metrics and a user study to provide a comprehensive analysis on the differences when annotating in synchronous and asynchronous settings. We first define and present several metrics for theme induction quality based on overlap, similarity/distance, association and verification that are well-established in the literature Ben-David and Ackerman (2008); Hoyle et al. (2021); Pacheco et al. (2023). These metrics are calculated using the sets of documents assigned to each theme by their respective system. While the Pacheco et al. (2023) system explicitly assigns a single theme for each document, the Fang et al. (2023) system provides a distribution of theme weights for each document. To ensure that results from both systems are comparable, we assign each document to the theme with the highest weight.
Finally, we summarize annotator experiences to provide a perspective on qualitative differences between the two settings. Combining dimensions in our evaluation methodology, we demonstrate that there are significant differences in consistency, cohesiveness, correctness, and user experiences between synchronous and asynchronous experiments.
5.1 Consistency
Coding individually always runs the risk that coders overgeneralize or miss key themes, resulting in “unconstrained or unsystematic” analysis (Cornish et al., 2014). We propose consistency as one dimension for measuring this systematicity. Building on the idea of inter-coder reliability, we define consistency as the measure of annotators eliciting the same themes from the same set of texts.
There are two obstacles to measuring annotation consistency when using semi-automated annotations. First, themes with similar names may not share the same annotated documents and themes with different names may share documents that are semantically similar, which we address by measuring the consistency of documents between given pairs of themes. Following the definition of consistency in Section 4 of Pacheco et al. (2023), we calculate the Jaccard similarity given as:
$J(A,B)=\frac{|A\cap B|}{|A|+|B|-|A\cap B|}$ (1)
where $A$ and $B$ are sets of documents for two themes with different coders. We then select the maximum Jaccard similarity for each (unique) theme, which we consider to be the “best theme match”, and calculate their averages for comparison across settings.
@@ 5.1 Consistency
Second, since the majority of documents are assigned a theme by the system, the same themes may contain a different assortment of documents but remain relatively semantically similar. We therefore also measure the semantic similarity between themes using their SBERT embeddings Reimers and Gurevych (2019). We do this in two ways: (1) Centroid similarity, where the similarity of two themes is the similarity between their two centroids, and centroids are calculated by taking the average of all document embeddings within the theme. (2) Group avg. similarity, where the similarity between two themes is the average of pair-wise similarity between all documents in the two themes. As in the Jaccard case, we report the maximum embedding similarity for each (unique) theme, which we consider to be the “best theme match”, and calculate their averages for comparison across settings. In all cases, we use the cosine similarity.
Table 1 shows the average maximum Jaccard and embedding similarities for each experiment. Comparing the results on the synchronous/asynchronous dimension, we find that synchronous groups produce themes with higher average consistency across the board. The centroid similarities and group average similarities in particular are statistically significant, supporting our hypothesis that different settings produce noticeably different coding results. These findings also demonstrate that the theoretical advantage for collaborative coding holds empirically true when using HitL systems as well: coders are more likely to miss themes or overgeneralize when working independently.
Comparing between consistency metrics, we find that centroid similarity is at or near statistical significance for all experiments and is much higher than the other metrics across the board. This finding is also in line with our expectations detailed above: since HitL systems automate a large portion of document assignments, small differences in user interactions can result in larger differences in document mappings. However, Jaccard similarity is still important to consider, especially for datasets comprised of documents that are very semantically similar.
5.2 Cohesiveness and Distinctiveness
Pacheco et al. (2023) Fang et al. (2023)
Intra-Theme Inter-Theme Intra-Theme Inter-Theme
Sync $\bf 0.51(0.08)^{*}$ $0.41(0.04)^{*}$ $0.52(0.10)$ $0.40(0.04)$
Async $0.45(0.10)^{*}$ $\bf 0.34(0.11)^{*}$ $0.52(0.10)$ $0.40(0.04)$
@@ 5.2 Cohesiveness and Distinctiveness
Table 2: Group Avg. Similarity within Annotator Groups (All data points). Themes are considered to be more cohesive if intra-theme similarity is high, and more distinctive if inter-theme similarity is low. Statistically significant using a two-sample unpaired t-test with $p<0.05$.
@@
A second dimension for determining the "systematicity, clarity, and transparency" of annotation outcomes is by evaluating the similarities and differences between themes within the same code-book (Cornish et al., 2014). We propose two metrics, cohesiveness and distinctiveness, to represent these similarities and differences respectively. Whereas consistency measures the systematicity and clarity of themes across annotators, cohesiveness and distinctiveness measure these qualities for themes by the same annotator. The distinctiveness of a theme is determined by comparing it to other themes created by the same annotator. On the other hand, the cohesiveness of a single theme is determined by its semantic homogeneity. A theme is said to be cohesive if its documents are similar to each other (measured by intra-theme similarity) and distinctive if it is dissimilar from documents in other themes within the same code-book (measured by inter-theme similarity). Intuitively, the purpose of grouping documents by theme is to create abstract representations of a dataset, where each theme represents a distinct facet of the data. If themes are not cohesive and distinctive, then it becomes hard to tell which theme a given document should belong to and the code-book falls apart.
@@
To evaluate these metrics for a single annotator (or annotator group), we calculate both the intra-theme similarity and the inter-theme similarity for theme in the code-book. Intra-theme similarity is calculated by taking the average of pair-wise similarity between all documents of the same theme. Inter-theme similarity for a given theme is calculated by taking the average pair-wise similarity of documents in that theme with documents in all other themes by the same annotator.
@@
A confounding factor in these measures is that both systems provide broad coverage of documents such that even distantly related documents may be assigned to a theme. In order to more accurately represent the cohesiveness and distinctiveness of themes in each experiment, we perform the same calculations on a subset comprised of only the top 25% of documents most closely related to each theme. For the Pacheco et al. (2023) system, this top quartile is selected using the distance from the centroid. For the Fang et al. (2023) system, the distribution of theme weights better reflect the assignments provided by the topic model and are used instead.
@@
Table 2 shows results for the whole dataset and Table 3 show the same metrics for the subset of the documents closest to each theme. Overall, we find that the intra-theme similarities are always higher than inter-theme similarities, which means that themes are at least moderately cohesive and distinctive across the board. For the Pacheco et al. (2023) experiments, we find that the difference between both types of metrics remains stable across collaboration modalities, and that similarities are higher in the synchronous setting. Our findings hint that themes may be more cohesive but not more distinctive in the synchronous setting.
Pacheco et al. (2023) Fang et al. (2023)
Intra-Theme Inter-Theme Intra-Theme Inter-theme
Sync $\bf 0.70(0.08)^{*}$ $0.52(0.07)^{*}$ $0.56(0.11)$ $0.39(0.05)$
Async $0.64(0.09)^{*}$ $\bf 0.46(0.13)^{*}$ $0.56(0.10)$ $0.39(0.05)$
Table 3: Group Avg. Similarity within Annotator Groups (25% most related). Themes are considered to be more cohesive if intra-theme similarity is high and more distinctive if inter-theme similarity is low. Statistically significant using a two-sample unpaired t-test with $p<0.05$.
The results in the Fang et al. (2023) experiments are nearly identical across collaboration modalities. Narrowing down to the top 25% of closest documents increases the difference between the two metrics, lining up with the intuitive notion that doing so increases both cohesiveness and distinctiveness. When comparing the two systems, we find that the themes coded using the Pacheco et al. (2023) system are significantly more cohesive when looking at the quartile subset, which matches our findings in Section 5.3.
5.3 Correctness
While the previous evaluations look at the code-book as a hole, we want to also evaluate how correct our document to theme assignments are. In order to verify this, we conducted a post-hoc analysis of how accurate theme assignments are on a held-out set of randomly sampled tweets, compared across synchronous and asynchronous settings. The data was sampled after the annotation process was over. This meant that for both systems, annotators had deemed the theme discovery and refinement to have reached saturation.
Following Section 5.2, we perform one analysis of correctness that considers the set of all tweets, as well as a second analysis considers only the top $25\%$ of tweets most closely related to each theme. For each analysis, we sampled $200$ tweets randomly and had two of the authors of the paper manually annotate whether the theme assigned to a tweet by the system makes sense or not. We did this for each system and collaboration setting (synchronous vs. asynchronous), with the exception of the Pacheco et al. (2023) synchronous setting, for which we report the results presented in the original paper. We made sure to follow the same sampling procedure so that results were comparable.
Results are presented in Tab. 4. We observe that for both systems, the accuracy yielded from synchronous annotations is higher than their asynchronous counter parts. We also see that as per our expectations, the top $25^{th}$ percentile has a higher accuracy than group average of all tweets.
However, it is interesting to observe the comparatively poor accuracy of the topic modeling system Fang et al. (2023) as compared to the theme discovery system Pacheco et al. (2023). We attribute this difference to the fact that topic modeling systems do not "learn" to classify themes, but rather learn the probability that a given word belongs to a particular topic. During evaluation, we found that the main problem arises when salient words are used in atypical contexts, or when trying to bifurcate terms that are used in opposing themes.
As an example, take the following headline regarding the COVID-19 vaccine mandate.
Indiana University students must comply with the school’s COVID-19 vaccine mandate for the fall semester after a federal judge refused on Monday to block the school’s inoculation requirement, one of the first rulings of its kind during the pandemic
The topic modeling system gave the highest wright to the topic Anger over mask mandates. But really, this is regarding vaccines. The term mandate biases the model to rank a topic about masks (the context that typically appears around the word "mandate") higher than a topic related to the COVID-19 vaccine, which this headline is about.
Our findings regarding the correctness can be summarized by acknowledging that synchronous teams of annotators achieve significantly higher assignment accuracy than asynchronous teams. We also find that topic modeling systems achieve un-competitive assignment accuracy versus theme discovery systems. Lastly, qualitative researchers may want to choose theme discovery tools if they want higher accuracy in document to theme assignments – but could opt for topic modeling systems if they want to put their focus on word usage patterns.
Pacheco et al. (2023) Fang et al. (2023)
All Top 25% All Top 25%
Sync ${}^{*}68\%$ ${}^{*}86\%$ $23\%$ $42\%$
Async $50\%$ $64\%$ $17\%$ $35\%$
Table 4: Correctness: Accuracy of Theme Assignment for Held-out Dataset. Results taken from Pacheco et al. (2023).
5.4 User Study
@@
We carried out semi-structured interviews with all the participants in the study after they were done with their coding task. We focused on trying to understand attitudes towards the task and tools, particularly with respect to the synchronous and asynchronous modalities. We summarize our main findings below.
Synchronous annotators found working in teams to ease the annotation process.
Both teams of annotators working synchronously using the Fang et al. (2023) described their perceived advantage of working in teams – broader perspectives by inviting view points from different annotators helped them contextualize diverse tweets easily. The annotators also valued the ability to collaboratively resolve contested tweets through discussions. Particularly, it was easy to "break ties" quickly through conversation when it came to mapping which words belonged to a given a topic. The annotators were also quicker to reach saturation, and they expressed that they thought this was possible because consolidation was possible on-the-fly. One of the annotators noted that a pre-requisite for successful synchronous coding stemmed from the rapport their annotation team shared, as they are all members of the same lab and were able to convey their viewpoints without hesitation.
Asynchronous annotators were more likely to notice inadequacies of the annotation tool they used.
While most of the highlights from the synchronous annotators focused on the process of annotation, the asynchronous annotators directed their feedback towards the way the tools functioned. Since the bulk of the annotations were carried in isolation, difficulties navigating the tool were more salient to them than to their synchronous counterparts. The synchronous annotators also met with similar issues, but were quick to solve it collaboratively. Even though all annotators (in both settings and systems) started with the same initial clusters of tweets, the synchronous annotators were able to find the operation they need performed quicker through in-team collaboration.
Annotators noted the lack of control and agency that the topic modeling system had over the theme discovery system.
@@
Annotators expressed loss of control and agency when using the topic modeling system Fang et al. (2023) over the theme discovery system Pacheco et al. (2023). One annotator noted that the controls to modify topics were not satisfactory – "the merge process did not offer the ideal amount of control and made it difficult to keep track of the theme groups".
Annotators felt that starting topics often negatively impacted their annotations.
@@
Since topic models start by a sampled cluster of data, both synchronous and asynchronous annotators were dissatisfied with some of the ways in which tweets belonging to different or opposite themes were placed under the same topic. One of the annotators commented – "Many Anti-Vax and Pro-Vax standpoints use the same words/phrases in their tweets, which the Fang et al. (2023) model groups together despite the stark difference in message between the two.". Since the topics contain a large number of documents, it is only feasible for annotators to include / exclude a very small subset as being part of the topic. Some users felt frustrated at being unable to achieve the purity of topic that they hoped for, often stopping when they found that none of their changes were affecting the topics as they intended. However, one annotator did note that the benefit of using a topic modeling system was that there was some amount of clustering already done - as opposed to the Pacheco et al. (2023) system where annotations had to be done from scratch, posing to be a small hurdle in the beginning.
@@ 6 Conclusion and Future Work
In this paper, we examined NLP-assisted qualitative research tools in different collaboration modalities. We conducted annotations on a vast dataset using two popular interactive coding tools, Pacheco et al. (2023) and Fang et al. (2023). We then designed an evaluation framework that describes the quality of the induced themes and their resulting document assignments under synchronous and asynchronous collaboration. We note that in general, the collaboration modality is a significant factor in determining if the quality of a system’s output. This is particularly true for theme discovery systems like Pacheco et al. (2023), where users have a wider bandwidth of interaction that can benefit from group consensus. We also observe that solutions based on topic modeling, although popular in the data analysis literature, can struggle with inducing cohesive themes and highly accurate document-level annotations.
While this study focuses on collaboration modalities, there are numerous other variables that can affect a tool’s efficacy in supporting qualitative coding. We believe that the evaluation framework that we proposed can be repurposed and expanded to evaluate a wide range of interventions, such as the underlying NLP technology, the interactive interface, the expertise and life experience of the coders, and the type of data being annotated. Through this paper, we hope to inform future work that gives way to more robust measures of an NLP tool’s capabilities as it pertains to qualitative research under realistic settings.

### v2

@@ 5 Evaluation
We use both descriptive metrics and a user study to provide a comprehensive analysis on the differences when annotating in synchronous and asynchronous settings. Our evaluation framework is comprised of three dimensions; consistency, cohesiveness & distinctiveness, and correctness, each of which uses metrics that are well-established in the literature Ben-David and Ackerman (2008); Hoyle et al. (2021); Pacheco et al. (2023).
Figure 2: Two sets of annotators use a particular HiTL system to find themes. Since the same theme can be named differently by different annotators, we find the best match. In this example, the annotator 1’s theme "VaxIsBad" has been matched with annotator 2’s theme "antivax". After aligning, we calculate the similarity between these two themes using methods like Jaccard Similarity or Centroid Distance.
Consistency. Coders always run the risk of overgeneralizing or missing key themes, resulting in unclear and unsystematic results (Cornish et al., 2014). We propose consistency as one dimension for measuring systematicity. Building on the idea of inter-coder reliability, we define consistency as the measure of annotators eliciting the same themes from the same set of texts (See Fig. 2). Once the dataset has been annotated using an interactive system, there are two obstacles to measuring consistency when using semi-automated annotations. First, themes with similar names may not share the same annotated documents. Similarly, themes with different names may share the same documents. We measure consistency using the overlap of documents between a given pairs of themes. To do this, we calculate the Jaccard similarity across the documents for two themes with different coders. We then select the maximum Jaccard similarity for each (unique) theme, which we consider to be the “best aligned theme”. In Fig. 2, this is shown in the middle where the colored lines indicate the best match obtained by maximum Jaccard overlap. After alignment, we calculate the average of all matched pairs for comparison across settings.
@@
We want to account for the possibility of themes containing a different assortment of documents that otherwise remain relatively semantically similar. To do this, we also measure the semantic similarity between themes using the S-BERT embeddings of the documents assigned to them Reimers and Gurevych (2019). We do this in two ways: (1) Centroid similarity, where the similarity of two themes is the similarity between their two centroids, and centroids are calculated by taking the average of all document embeddings within the theme. (2) Group avg. similarity, where the similarity between two themes is the average of pair-wise cosine similarity between all documents in the two themes. As in the Jaccard metric, we report the maximum embedding similarity for each (unique) theme, which we consider to be the “best theme match”, and calculate their averages for comparison across settings.
Table 1 shows the average maximum Jaccard and embedding similarities for each experiment. We find that synchronous groups produce themes with higher average consistency for the topic modeling and relational systems. For the LLM-based system, we did not observe statistical significant differences across collaboration settings. Interestingly, the LLM-based system offered the least amount of intervention/refinement operations for users during the coding process. Whereas the other systems allowed for operations such as splitting and merging topics (topic modeling), and defining relations between concepts (relational approach), the LLM system only allowed users to alter the definition of the theme provided in the prompt. We hypothesize that the more interventions users can make, the more advantage they can take of the deliberation afforded by synchronous coding.
Topic Model Relational LLM-based
Intra-Theme Inter-Theme Intra-Theme Inter-Theme Intra-Theme Inter-Theme
All Sync $0.52(0.10)$ $0.40(0.04)$ $\bf 0.51(0.08)^{*}$ $0.42(0.05)^{*}$ $\bf 0.44(0.06)$ $0.40(0.04)$
Async $0.52(0.10)$ $0.40(0.04)$ $0.45(0.10)^{*}$ $\bf 0.34(0.11)^{*}$ $0.43(0.05)$ $\bf 0.39(0.04)$
Top 25% Sync $0.56(0.11)$ $0.39(0.05)$ $\bf 0.70(0.09)^{*}$ $0.52(0.07)^{*}$ $0.63(0.07)$ $0.55(0.05)$
Async $0.56(0.11)$ $0.39(0.05)$ $0.64(0.09)^{*}$ $\bf 0.46(0.13)^{*}$ $0.63(0.05)$ $\bf 0.54(0.05)$
@@
Table 2: Group Avg. Similarity within Annotator Groups. Themes are considered to be more cohesive if intra-theme similarity is high and more distinctive if inter-theme similarity is low. Statistically significant using a two-sample unpaired t-test with $p<0.05$.
Figure 3: Once an annotator has identified themes and they have been propagated the full dataset, we calculate intra-theme similarity by measuring the avg. of the pairwise distances between each document within a theme (left). We calculate inter-theme similarity by measuring the avg. of pairwise distances between each document in a theme and documents assigned to all other themes (right)
@@
Cohesiveness and Distinctiveness. Another dimension for determining the systematicity and clarity of annotation outcomes is by evaluating the similarities and differences between themes within the same code-book. We propose two metrics to measure this: cohesiveness and distinctiveness. A theme is said to be cohesive if its documents are similar to each other (measured by intra-theme similarity) and distinctive if it is dissimilar from documents in other themes within the same code-book (measured by inter-theme similarity). Intuitively, the purpose of grouping documents by theme is to create abstract representations of a dataset, where each theme represents a distinct facet of the data. If themes are not cohesive and distinctive, then it becomes hard to tell which theme a given document should belong to and the code-book falls apart.
@@
Figure 3 shows how to evaluate these metrics for a single annotator (or annotator group). We calculate both the intra-theme similarity and the inter-theme similarity for all the themes in the code-book. Intra-theme similarity is calculated by taking the average of pair-wise similarity between all documents of the same theme. Inter-theme similarity for a given theme is calculated by taking the average pair-wise similarity of documents in that theme with documents in all other themes.
@@
A confounding factor in these measures is that all systems provide broad coverage of documents such that even distantly related documents may be assigned to a theme. To more accurately represent the cohesiveness and distinctiveness of themes in each experiment, we perform the same calculations on a subset comprised of only the top 25% of documents most closely related to each theme. For the relational and LLM-based systems, this top quartile is selected using the distance from the centroid. For the interactive topic model, we use the weights assigned by the model.
@@
Table 2 shows results for both the whole dataset as well as the subset of the documents closest to each theme. Overall, we find that the intra-theme similarities are always higher than inter-theme similarities, which means that themes are at least moderately cohesive and distinctive across the board. For the relational system experiments, we find that the difference between both types of metrics remains stable across collaboration modalities, and that similarities are higher in the synchronous setting. Our findings suggest that themes may be more cohesive but not more distinctive in the synchronous setting.
Results for the interactive topic model and LLM-based systems are nearly identical across collaboration modalities. Given that the operations to characterize themes offered by these two systems are substantially limited (i.e., adding and removing words for the topic model, and refining theme definitions for the LLM system), these results are not surprising. Again, we see that the wider the range of operations afforded to characterize a theme, the more advantage can be taken from synchronous deliberation.
Correctness. Interactive systems allow users to automate large portions of the coding process at the risk of producing inaccurate theme assignments. To estimate how correct the outputs of each system are, we conduct a post-hoc analysis by manually checking a randomly selected sample of $1,200$ document-theme pairs ($200$ per experimental setting). To ensure that our sample is representative of the overall dataset for each experimental setting, we split the data into quartiles based on document relatedness to each theme and select a uniform sample of themes and relatedness scores. As in the previous section, relatedness is calculated using the theme weight distribution for the interactive topic model and distance from the centroid for the the other two systems. To assess reliability, each assignment is evaluated by two annotators, with a third weighing in for tie breaks. We also find that human evaluators have moderate-to-high agreement when assessing system outputs (with an overall Krippendorff’s $\alpha$ of $0.632$), suggesting that we can trust these estimations.
Figure 4 shows the correctness results for each quartile sample per experiment. First, we observe that the relational system is not only the most accurate, but it shows negligible correctness differences in synchronous vs. asynchronous configurations. This an encouraging result, given that this system took the most advantage of synchronous deliberation based on our other metrics of quality. The other two systems showed marked differences, with the topic modeling approach producing more accurate assignments in asynchronous operation, and the LLM producing more accurate assignment in a synchronous paradigm. This observation leads us to believe that interactive systems that introduce less and weaker inductive biases from interaction are more sensitive to choices of configuration – and researchers should be aware of these possibilities when designing experiments. Second, we find that the correctness of different approaches tapers differently, based on the distance of the example from the theme. In other words, as the data samples resemble less like the human labeled ones, some models handle it more robustly than others. In our observation, relational approaches and LLM methods outperform topic models when it comes to assigning themes to distant examples.
Figure 4: Correctness w.r.t. distance from theme.
6 User Study and Recommendations
@@
We carried out semi-structured interviews with the participants in the study. We focused on trying to understand attitudes towards the task and tools, particularly with respect to the synchronous and asynchronous modalities. Interview questions can be found in App. C. We summarize our main findings below.
Participants found working in teams eases the coding process. Teams of annotators working synchronously described their perceived advantage of working in teams – broader perspectives helped them contextualize diverse tweets easily. The annotators also valued the ability to collaboratively resolve disagreements through discussions. They cited it being easy to "break ties" quickly through conversation. In addition to easing the user experience, we also note that the consistency and cohesiveness of annotations is higher in synchronous settings. Based on this evidence, future work should consider designing systems that explicitly support synchronous deliberation.
Asynchronous coders were more likely to notice inadequacies of the tool they used. Asynchronous annotators directed most of their feedback towards the way the tools functioned. Since the bulk of the annotations were carried in isolation, difficulties navigating the tool were more salient to them than to their synchronous counterparts. Based on these findings, tool designers should focus on improving the asynchronous coding experience.
@@
Annotators noted a lack of control and agency in the topic modeling system. Annotators expressed loss of control and agency when using the topic modeling system Fang et al. (2023) over other systems. One annotator noted that the controls to modify topics were not satisfactory – "the merge process did not offer the ideal amount of control and made it difficult to keep track of the theme groups". Based on these findings, tool designers should consider abstracting users away from the intricacies of the underlying NLP method.
@@
Annotators felt that starting topics often negatively impacted their annotations. Both synchronous and asynchronous annotators were dissatisfied with some of the ways in which tweets belonging to different or opposite themes were placed under the same topic in the topic modeling system. In contrast to the clustering approach used by the other two systems, which appeared to yield cleaner partitions. One of the annotators commented – "Many Anti-Vax and Pro-Vax standpoints use the same words/phrases in their tweets, which the Fang et al. (2023) model groups together despite the stark difference in message between the two.". Some users felt frustrated at being unable to achieve the purity of topic that they hoped for, often stopping when they found that none of their changes were affecting the topics as they intended. However, one annotator did note that the benefit of using a topic modeling system was that there was some amount of theme induction done already - as opposed to the other two systems, where clusters did not have any representative words attached to them. Based on these findings, researchers should investigate ways to effect more change during interactions for topic modeling systems, as well as consider generating some initial explanations to characterize clusters for alternative approaches.
Using LLMs to conduct thematic coding at scale is expensive and less reliable. Despite the recent advances in the reasoning capabilities of LLMs, we observe that their performance at scale is not effective enough to be used for the purpose of propagating annotations. Additionally, running LLMs for the $5$ annotation configurations came with significant time and resource costs. This opens an opportunity for NLP researchers to make LLMs more reliable inductive reasoners, and to come up with prompting strategies that can allow LLMs to reliably classify documents in bulk, especially when working at scale.
Future work developing interactive systems should take into account the amount of control the system offers to the user. Quantitatively, we find that the topic modeling system produced results that were less correct across the board. This matches with our user study where users were unable to disentangle themes that had a high degree of lexical similarity. Despite being able to refine the topic model such that the highest weighted documents matched their defined themes, users had little control over whether or not the rest of the documents in the dataset were faithfully assigned.
@@ 7 Conclusion and Future Work
In this paper, we examined three categories of NLP-assisted qualitative research tools in different collaboration modalities, and conducted inductive and deductive coding on a large dataset of English texts. We designed an evaluation framework that describes the quality of the induced themes and their resulting document assignments under synchronous and asynchronous collaboration. We note that in general, the collaboration modality is a significant factor in determining if the quality of a system’s output. This is particularly true for systems like Pacheco et al. (2023), where users have a wider bandwidth of interaction that can benefit from group consensus, and provide strong inductive biases for deductive coding. We also observe that solutions based on topic modeling, although popular in the data analysis literature, can struggle with inducing cohesive themes and accurate code assignments. Finally, we show that while LLM-based solutions show promise, they pose significant challenges when it comes to coding at scale.
While this study focuses on collaboration modalities, there are numerous other variables that can affect a tool’s efficacy for qualitative coding. We believe that our proposed evaluation framework can be repurposed and expanded to evaluate a wide range of interventions, such as the underlying NLP technology, the interactive interface, the expertise and life experience of the coders, and the type of data being annotated. Through this paper, we hope to inform future work that gives way to more robust measures of an NLP tool’s capabilities as it pertains to qualitative research under realistic settings.